OpenStudy (anonymous):

write in logarithmic form (2/3)^-1 = 3/2

OpenStudy (ghazi):

on taking log both the sides you will get \[\log ((2/3)^ {-1})=\log (3/2)\] which gives you\[-1*\log (2/3)=\log(3/2)\]

OpenStudy (ghazi):

\[\log a -\log b= \log (a/b)\]

OpenStudy (ghazi):

can you do the rest?

OpenStudy (ghazi):

log (3/2) + log (2/3) and Log a + log b = Log (a*b)

OpenStudy (anonymous):

i think by "logarithmic form" they mean change \[b^x=y\] to \[log_b(y)=x\]

OpenStudy (anonymous):

so \((\frac{2}{3})^{-1}=\frac{3}{2}\) becomes \[\log_{\frac{2}{3}}(\frac{3}{2})=-1\]

OpenStudy (ghazi):

but what i have done, is also one of the possibilities

OpenStudy (anonymous):

uh... i stuck with -log(2/3)=log(3/2) because when i do evaluate that i get 1.

OpenStudy (ghazi):

you will get zero log a + log b = log (a*b)= log ( 2/3 * 3/2) = log 1 =0